# Efficient and Generalizable Certified Unlearning: A Hessian-free Recollection Approach

### v3

@@ 4.2 Generalization Properties Analysis
$F(\tilde{\mathbf{w}}_{E,B}^{-U})-\mathbb{E}[F\left(\mathbf{w}^{*}\right)]=\mathcal{O}\Big{(}\frac{4L^{2}}{\lambda(n-m)}+\rho^{T}\frac{2L^{2}}{\lambda}+\frac{2L\eta C\sqrt{d}\sqrt{\ln{(1/\delta)}}}{\epsilon}\zeta_{T}^{-U}\!\Big{)}.$ (14)
The proof is included in Subsection B.6, where we perform an asymptotic analysis. Specifically, the result of our proposed method is expected to converge to $\mathcal{O}(\frac{L^{2}}{\lambda(n-m)})$ as $T$ approaches infinity. It is noteworthy that, under these conditions and the fact that $n=\Omega(m)$, our proposed approximator outperforms the performance of previous methods [2, 3, 4] which are greater than $\mathcal{O}(\frac{L^{2}}{\lambda(n-m)})$. Furthermore, to measure the maximum number of samples to be forgotten while still ensuring good excess risk guarantees, we analyze the deletion capacity of our proposed scheme, similar as in [2]. Due to the space limit, we defer the definition and analysis to Subsection E.1, where we compare our main theoretical results for generalization and deletion capacity with previous works.
4.3 Efficient Unlearning Algorithm Design
Based on the above methods and the analysis of Theorem 4, we now propose a memory and computationally efficient unlearning scheme in Algorithm 1. To our best knowledge, our method is currently one of the most efficient, as it only requires a single vector addition by utilizing offline computed statistics. Besides, our proposed algorithm also ensures complete forgetting by deleting the approximator of the corresponding sample post-unlearning, rather than storing such statistics.
We present our approach in Algorithm 1. In Stage I, we bound approximation error through gradient clipping and perform model training normally. Upon completing the model training, we proceed to Stage II, where we compute the unlearning statistics vector $\mathbf{a}_{E,B}^{-u}$ for each sample $u$. To reduce computational complexity, we leverage the HVP technique. The computed approximators $\{\mathbf{a}_{E,B}^{-u_{j}}\}_{j=1}^{n}$ are then stored for further use. A deletion request initiates Stage III. In this stage, the indistinguishability of the $(\epsilon,\delta)$-unlearning guarantee, as defined in Definition 1, is achieved through a single vector addition with injected noise. Notably, our unlearning approach does not necessitate accessing any datasets during Stage III. Moreover, in contrast to prior works that require retaining privacy-sensitive statistics to handle subsequent forgetting requests, our approach eliminates the need to retain statistics post-unlearning by simply discarding them. Lastly, our method necessitates minimal modifications to existing training frameworks. This characteristic renders our approach orthogonal to any training innovations and opens up possibilities for its future expansion, e.g., our method can be combined with fine-tuning to enhance performance. Specifically, this involves training on a small number of epochs with the remaining dataset, and then adding the corresponding approximators during this period to the original approximators. The detailed strategy is described in Subsection E.3.
1 Stage I: Learning model $\mathbf{w}_{E,B}$ on dataset $S=\{z_{i}\}_{i=1}^{n}$:
2 for  $e=0,1\cdots,E$ do
@@ 1 Stage I: Learning model $\mathbf{w}_{E,B}$ on dataset $S=\{z_{i}\}_{i=1}^{n}$:
4             Compute gradient: $\mathbf{g}\!\leftarrow\!\frac{1}{|\mathcal{B}_{e,b}|}\sum_{i\in\mathcal{B}_{e,b}}\!\!\!\nabla\ell\left(\mathbf{w}_{e,b};z_{i}\right)\times\max(1,\frac{C}{\left\|{\nabla\ell\left(\mathbf{w}_{e,b};z_{i}\right)}\right\|})$,
@@ 5             Gradient descent: $\mathbf{w}_{e,b+1}\leftarrow\mathbf{w}_{e,b}-{\eta_{e,b}}{\mathbf{g}}$
9
10Stage II: Pre-computing and Pre-storing statistics $\mathcal{T}(\mathcal{S})=\{\mathbf{a}_{E,B}^{-u_{j}}\}_{j=1}^{n}$:
11 for  $j=1,2\cdots,n$ do
12       Recursive computation by using HVP: $\mathbf{a}_{E,B}^{-u_{j}}\leftarrow\sum_{e=1}^{E}\mathbf{M}_{e,b(u_{j})}\nabla\ell(\mathbf{w}_{e,b(u_{j})};u_{j})$,
13       where $\mathbf{M}_{e,b(u_{j})}=\frac{\eta_{e,b(u_{j})}}{|\mathcal{B}_{e,b(u_{j})}|}\prod^{E}_{k=e}\prod_{b=b(u_{j})+1}^{B-1}\!(\mathbf{I}-\frac{\eta_{k,b}}{|\mathcal{B}_{k,b}|}\mathbf{H}_{k,b})$
@@
15Stage III: Unlearning when user requests to forget the subset $U$=$\{u_{j}\}_{j=1}^{m}$ on model $\mathbf{w}_{E,B}$:
16 Compute: $\bar{\mathbf{w}}_{E,B}^{-U}\!\leftarrow\!\mathbf{w}_{E,B}\!+\!\mathbf{a}_{E,B}^{-U},$ $c\!\leftarrow\!{2\eta C}\frac{\sqrt{2\!\ln\!{(1.25/\delta})}}{\epsilon}{\zeta_{T}^{-U}}\!\!$    Delete statistics: $\{\mathbf{a}_{E,B}^{-u_{j}}\}_{j=1}^{m}$
Sample: $\sigma\sim\mathcal{N}(0,c\mathbf{I})$, Return: $\tilde{\mathbf{w}}_{E,B}^{-U}=\bar{\mathbf{w}}_{E,B}^{-U}+\sigma$
Algorithm 1 Hessian-free Online Unlearning (HF) Algorithm
4.4 Comparison to Existing Approaches
We provide a brief overview of previous second-order unlearning works and compare the complexity of our method with these works in terms of pre-computation, storage, and unlearning time.
Newton Step (NS). Sekhari et al. [2] proposed an efficient certified unlearning method by pre-storing the Hessian $\sum_{i\!=\!1}^{n}\!\!\nabla^{2}\ell(\mathbf{w}_{E\!,B};\!z_{i})$. When a request arrives to forget $m$ samples $U$, we obtain an unlearned model $\mathbf{w}^{-U}_{E,B}$ by adding the approximator $\mathbf{a}^{-U}_{\text{NS}}$ to original learned model $\mathbf{w}_{E,B}$.
$\!\!\!\mathbf{a}^{-U}_{\text{NS}}\!\!=\!\frac{1}{n\!-\!m}\mathbf{H}_{\text{NS}}^{-1}\sum_{i\in U}\!\nabla\ell(\mathbf{w}_{E,B};z_{i}),\mathbf{H}_{\text{NS}}\!=\!\frac{1}{n\!-\!m}\Big{(}\!\sum_{i=1}^{n}\!\nabla^{2}\ell(\!\mathbf{w}_{E,B};z_{i})\!-\!\sum_{i\in U}\nabla^{2}\ell(\mathbf{w}_{E,B};z_{i}\!)\!\Big{)}.$ (15)
Infinitesimal Jacknife (IJ). Suriyakumar et al. [3] extends the work of NS by employing proximal IJ, and the main difference lies in the strategy of pre-computing and storing the inverse Hessian $\mathbf{H}_{\text{IJ}}$. It enables forgetting by adding the approximator $\mathbf{a}^{-u}_{\text{IJ}}$ to the learned model $\mathbf{w}_{E,B}$.
$\mathbf{a}^{-u}_{\text{IJ}}$ $=\frac{1}{n}\mathbf{H}_{\text{IJ}}^{-1}\nabla\ell(\mathbf{w}_{E,B};u),\ \mathbf{H}_{\text{IJ}}$ $=\frac{1}{n}\sum_{i=1}^{n}\nabla^{2}\ell(\mathbf{w}_{E,B};z_{i}).$ (16)
Comparison of Pre-Computation. First, for a single data point $u$, Algorithm 1 entails computing the HVP of the gradient vectors $\nabla\ell(\mathbf{w}_{e,b};u)$ $T$ times, which requires $\mathcal{O}(Td)$ running time. To compute the our proposed hessian-free (HF) approximators $\mathbf{a}^{-u}_{\text{HF}}$ for all $n$ samples, the complexity of running time is $\mathcal{O}(nTd)$. Our proposed method is highly efficient when dealing with overparameterized deep models, especially when $d\gg n$. This is because computing the full Hessian for NS requires $\mathcal{O}(nd^{2})$ running time, and computing and inverting the full Hessian $\mathbf{H}^{-1}_{\text{IJ}}$ for IJ requires $\mathcal{O}(d^{3}+nd^{2})$.
Comparison of Storage. Algorithm 1 involves storing a vector of model parameter size for $n$ data, requiring the storage of $\mathcal{O}(nd)$. This is lower than the overhead introduced by storing matrices for NS and IJ, which is $\mathcal{O}(d^{2})$, when the dataset size $n$ is lower than the dimension $d$ of model parameters [^2] It’s noteworthy that in the realm of deep neural networks for most computer vision tasks, dataset size $n$ usually falls short of the model parameter dimension $d$. Even for neural language models, the empirical data-to-parameter scaling law is expected to remain at $n\sim d^{0.74}$, as suggested in [31]..
Comparison of Unlearning Computation. When a forgetting request with $m$ data to be forgotten arrives, Algorithm 1 utilizes precomputed and prestored approximators to achieve forgetting, requiring a simple vector addition that takes $\mathcal{O}(md)$ time. For NS, it requires computing and inverting a different Hessian that depends on the user requesting the deletion. This necessitates substantial computational overhead of $\mathcal{O}(d^{3}+md^{2}+md)$ each time a deletion request arrives. Although IJ reduces this to $\mathcal{O}(md^{2}+md)$, it is still not as efficient compared to our method.
5 Experiments
We devised the following experimental components: verification and application. Verification experiments are centered on evaluating the disparity between the unlearned and retrained models, across both convex and non-convex settings. Application experiments are geared towards evaluating the performance of different unlearning algorithms with a specific focus on determining the performance, pre-computation/storage, and unlearning time of these methods. Considering the space limit, we present a portion of the results from the verification experiments and application experiments, while deferring the remaining experiments to the Appendix C.
5.1 Verification Experiments
Our principal aim is to validate the differences between the proposed Hessian-free (HF) approximator $\mathbf{a}^{-U}_{\text{HF}}$ and the actual $\mathbf{w}_{E,B}^{-U}\!-\!\mathbf{w}_{E,B}$, where a smaller difference indicates the unlearned model is closer to retrained model and also reflects the degree of data forgetting. Therefore, we evaluate the degree of forgetting from two perspectives: distance and correlation. Specifically, we use the $L_{2}$ norm $\|\mathbf{w}_{E,B}^{-U}\!\!-\!\mathbf{w}_{E,B}\!\!-\!\mathbf{a}^{\!-U}\|$ to measure the distance metric. A smaller distance metric indicates the approximators accurately capture the disparity between the retrained and learned models. Additionally, we utilize the Pearson [32] and Spearman [33] coefficient to assess the correlation between $\mathbf{a}^{-U}$ and $\mathbf{w}_{\!E,B}^{-U}\!-\!\mathbf{w}_{\!E,B}$ by mapping them from high-dimensional space to scalar loss values, i.e. the approximate loss change $\ell(\mathbf{w}_{E,B}\!+\!\mathbf{a}^{\!-U}\!;U)\!-\!\ell(\mathbf{w}_{E,B};U)$ and the actual loss change $\ell(\mathbf{w}^{-U}_{E,B};U)\!-\!\ell(\mathbf{w}_{E,B};U)$ on the forgetting dataset $U$. The resulting correlation scores range from -1 to 1, where the higher the correlation score, the more the unlearned model performs like the retrained model.
Configurations: We conduct experiments in both convex and non-convex scenarios. Specifically, we trained a multinomial Logistic Regression (LR) with total parameters $d=7850$ and a simple convolutional neural network (CNN) with total parameters $d=21840$ trained on MNIST [34] for handwriting digit classification. We apply a cross-entropy loss function and the inclusion of an $L_{2}$ regularization coefficient of $10^{-6}$ to ensure that the loss function of LR is strongly convex. For LR, training was performed for 100 epochs with stepsize of 0.05 and full batch. For CNN, training was carried out for 20 epochs with stepsize of 0.05 and a batch size of 64. Given these configurations mentioned above, we separately assessed the distance and correlation between our approximators $\mathbf{a}^{-U}_{\text{HF}}$, $\mathbf{a}^{-U}_{\text{NS}}$, $\mathbf{a}^{-U}_{\text{IJ}}$ at deletion rates in the set {1%, 5%, 10%, 15%, 20%, 25%, 30%}. Following the suggestion in [35], a damping factor of 0.01 is added to the Hessian to ensure its invertibility when implementing NS and IJ. Besides, we conducted experiments with 7 random seeds to obtain average results. For the sake of clear visualizations, we use the minimum and maximum values as error bars.
The convex setting. As shown in Figure 1 (a), our approach outperforms previous works. First, by distance evaluation, the approximation error $\|\Delta_{E,B}^{-U}\|$ of the proposed method is lower than that of the previous NS and IJ works. At a deletion rate of 30%, the approximation error is 0.171638, slightly lower than that of IJ and NS, which have errors of 0.178244 and 0.178246, respectively. Second, in assessing the loss change for the forgotten dataset $U$, our proposed method more accurately captures the changes in actual $\ell(\mathbf{w}^{-U}_{E,B};U)-\ell(\mathbf{w}_{E,B};U)$. In particular, when removing 30% of the samples, the proposed method maintains a high correlation, with Pearson and Spearman being 0.96 and 0.95, respectively. These evaluations demonstrate the efficacy of both Hessian-free and Hessian-based methods in accurately approximating the retrained model within convex settings. Our proposed approach not only achieves a closer approximation to the retrained model but also exhibits lower complexity compared to previous methods. This is substantiated by the results of Application Experiments I, which provide empirical evidence supporting our theoretical findings.
The non-convex setting. As illustrated in Figure 1 (b), proposed HF method demonstrates superior performance, exhibiting less dependency on random variations compared to NS and IJ. When 30% of the samples are removed, our method maintains a lower approximation error of 0.96, surpassing the performance of IJ and NS. Our methods achieve Spearman and Pearson correlation coefficient of 0.80 and 0.74, respectively, accurately capturing actual loss changes, in contrast to IJ and NS which are far from the actual values in non-convex settings. In Subsection C.1, we explain the volatility of the results in the prior works on non-convex setting, and provide more large-scale results, which is not achievable in NS and IJ [2, 3, 4] for their high complexity requirements and restrictive assumptions.
(a) Convex Setting
(b) Non-Convex Setting
Figure 1: Verification experiments I on (a) LR and (b) CNN, respectively. The left of (a)(b) shows the distance to the retrained model, i.e. $\|\mathbf{w}_{E,B}^{-U}\!-\!\mathbf{w}_{E,B}\!\!-\!\mathbf{a}^{-U}\|$. The right of (a)(b) shows the correlation between the approximate and actual loss change on the forgetting data.
5.2 Application Experiments
Furthermore, we compare the cost and performance of our method with the previous certified unlearning methods. Our evaluations are conducted from two perspectives: runtime and utility. Specifically, run-time focuses on the time spent precomputing unlearning statistics and the speedup of the unlearning algorithm compared to the retraining algorithm. Moreover, we evaluate the utility by the test accuracy of the unlearned model to ensure that generalization performance is not compromised. We train LR and simple CNN on MNIST with 20% data to be forgotten, which is identical to the aforementioned verification experiments I.
Table 2: Application experiments I. An online scenario with 20% data to be forgotten, where each unlearning request involves forgetting a single data point with each execution. The values in parentheses represent the difference in test accuracy compared to the retrained model. We compute the total computational/storage cost for all samples and test the accuracy of the final unlearned model.
Model Unlearning Computation PreComputaion Storage Test Accuracy (%)
Runtime (Sec) Speedup Runtime (Sec) (GB) Unlearned model
NS Logistic 5.14×$\times 10^{2}$ 2.39$\times$ 2.57×$\times 10^{3}$ 0.23 87.50 (-0.75)
CNN 5.82×$\times 10^{3}$ 0.05$\times$ 2.91×$\times 10^{4}$ 1.78 83.50 (-10.25)
IJ Logistic 1.84×$\times 10^{0}$ 665$\times$ 2.57×$\times 10^{3}$ 0.25 87.50 (-0.75)
CNN 6.81×$\times 10^{0}$ 45$\times$ 2.91×$\times 10^{4}$ 1.78 82.75 (-11.00)
HF Logistic 2.74×$\times 10^{-3}$ 448,060$\times$ 1.95×$\times 10^{1}$ 0.03 87.75 (-0.50)
CNN 3.23×$\times 10^{-2}$ 9,940$\times$ 5.34×$\times 10^{2}$ 0.08 91.50 (-2.25)
Application experiment results. Table 2 demonstrates that the proposed algorithm outperforms other baselines by a significant margin. Strikingly, the proposed algorithm outperforms benchmarks of certified unlearning regarding all metrics with millisecond-level unlearning and minimal performance degradation in both convex and non-convex settings. This demonstrates that the Hessian-free perspective method exhibits great potential for over-parameterized models compared to previous Hessian-based approaches. In Subsection C.2, we provide larger scale results, which is not achievable in previous work [2, 3, 4] due to their high complexity requirements and the restrictive assumptions.
6 Conclusion
In this work, we proposed a novel Hessian-free certified unlearning method. Specifically, we introduced an approximator based on affine stochastic recursion to characterize the trajectory discrepancy between the learned and retrained models. Our analysis does not involve inverting Hessian and thus can be extended to non-convex non-smooth objectives. Additionally, we provide theoretical results of generalization, unlearning guarantees, deletion capacity, and computation/storage complexity that are superior to state-of-the-art methods. Based on this analysis, we developed a certified unlearning algorithm that can achieve low-complexity and near-instantaneous unlearning. We employ HVP and gradient clipping techniques to mitigate the computation time and bound approximation error, respectively. We showed that proposed method can reduce the precomputation time, storage, and unlearning time (per sample) to $\mathcal{O}(nd)$, $\mathcal{O}(nd)$, $\mathcal{O}(d)$, respectively. Experimental results validate that our work benefits certified unlearning with reduced computation/storage and improved accuracy.
Appendix A An Expanded Version of The Related Work
Machine Unlearning, the terminology can be traced back to the paper of Cao et al. [8], and their work defined the notion of "the completeness of forgetting". There are two main categories based on the completeness of forgetting: Exact Unlearning and Approximate Unlearning.
Exact Unlearning aims to accelerate the speed of retraining and completely eliminate the influence of forgotten data on learned model. The majority of existing Exact Unlearning methods are derived from the work of Bourtoule et al. [9], which proposed the notable SISA framework, grounded in ensemble learning techniques. Specifically, the training dataset is partitioned into mutually disjoint shards, and then sub-models are trained on these shards. Upon unlearning, SISA only needs to retrain the sub-model correlated with the relevant data shard and then make the final prediction by assembling the knowledge of each sub-model, which significantly lowers retraining computational costs. Under the SISA framework, other Exact Unlearning methods have also been proposed, e.g., GraphEraser [13] extended SISA to graph learning by partitioning the original training graph into disjoint shards, parallelly training a set of shard models, and learning an optimal importance score for each shard model. RecEraser [14] extended SISA to recommendation systems. KNOT [36] adopted the SISA for client-level asynchronous federated unlearning during training and introduced a clustered mechanism that divides clients into multiple clusters. Other methods, such as DaRE [10] and HedgeCut [11], addressed the issue that SISA cannot be applied to decision trees and proposed exact unlearning solutions; and [37] aimed to achieve forgetting in K-means clustering. Furthermore, ARCANE [12] introduced a class-based data partitioning strategy that enhances the performance of SISA.
However, the SISA framework adopts a structure that completely separates the data, resembling ensemble learning approaches where predictions or models are aggregated only at the end. This separation restricts the model training process from fully leveraging the entirety of the segmented datasets and also hampers its extensibility, consequently leading to lower accuracy and performance compared to a model trained on the entire unsegmented dataset. In an experimental evaluation conducted by Koch et al. [38], it was found that simply down-sampling the dataset to achieve a more balanced single shard outperforms SISA in terms of both unlearning speed and performance.
Approximate Unlearning thus seeks to minimize the impact of forgetting data to an acceptable level to trade off for computational efficiency, reduced storage costs, and flexibility. There are many approaches to realizing Approximate Unlearning, such as [39] developed a distillation-based method; [40] adopted a gradient ascent approach on the forgotten data to achieve forgetting. Among these, second-order methods have been widely studied, with most of these methods (e.g., [2, 3, 4, 6, 26, 27]) inspired by influence function studies [41, 35, 42], aiming to extract information from the Hessian of the loss function and achieve forgetting through limited model updates. In particular, inspired by differential privacy [43], Guo et al. [1] introduced a $(\epsilon,\delta)$-certified unlearning definition, anticipating ensuring the output distribution of the unlearning algorithm is indistinguishable from that of retraining. Furthermore, due to the high computational cost of the Hessian, Fisher-based methods [17, 44, 45] have emerged to approximate the Hessian. Our paper focuses on Hessian-based certified unlearning.
Among these Hessian-based certified unlearning methods, Sekhari et al. [2]; Suriyakumar et al. [3]; Liu et al. [4] established the theoretical foundation for Hessian-based methods and provided the generalization result in terms of the population risk, as well as derived the deletion capacity guarantee. In the definition of certified unlearning, [2, 3, 4] align with our method, which lies in pre-computing statistics of data that extract second-order information from the Hessian and using these ‘recollections’ to forget data. In particular, Sekhari et al. [2] pointed out the importance of providing the generalization performance after machine unlearning and proposed achieving forgetting upon the arrival of a forgetting request using a Newton step through the precomputation and storage of the Hessian offline. Suriyakumar et al. [3, 4] extended the work of Sekhari et al. [2], attaining strong convexity through a non-smooth regularizer by leveraging the proximal IJ. Moreover, by trading off the offline computation time to invert the Hessian, the unlearning time was reduced to $\mathcal{O}(d^{2})$. Liu et al. [4] further developed an $(\epsilon,\delta)$-certified unlearning algorithm for minimax models.
However, these second-order certified unlearning methods have encountered persistent challenges that are difficult to overcome. That is, existing certified unlearning methods, especially those focused on establishing theoretical results, have largely focused on strongly convex and smooth objectives. Moreover, prior second-order methods require expensive Hessian computations, thus suffering from the curse of model parameter dimensionality, making it challenging to apply to most deep model.
Appendix B Proofs
B.1 Detailed Analysis of Eq. 10
Here, we provide a more detailed proof and explanation for (10). Specifically, we first consider the impact of sample $u$ during the $e$-th epoch, and then extend the analysis to all epochs.
It should be noted that a necessary setting for our method is that the retraining and learning process have the same initialization, model architecture, and training process.
Analysis of (10).
Since each sample participates in updates only once per epoch, without loss of generality, for the $e$-th epoch, we define the sample $u$ to be involved (sampled) in the $b(u)$-th update of the $e$-th epoch. Therefore, for a learning model in (3) and a retraining model obtained without the knowledge of $u$ in (4), we have that the following one step updates,

### v4

@@ B.2 Deletion Capacity Guarantee
$m_{\epsilon,\delta,\gamma}^{\Omega,\bar{\Omega}}(d,n)$ $:=\max\left\{m\mid\mathbb{E}\left[\max_{U\subseteq\mathcal{S}:|U|\leq m}F(\bar{\Omega}(\Omega(\mathcal{S});T(\mathcal{S}))-F(\mathbf{w}^{*})\right]\leq\gamma\right\},$ (17)
Similar to Sekhari et al. (2021), we provide upper bound and lower bound of deletion capacity for our algorithm. We first give the upper bound of deletion capacity.
Theorem 7 (Upper Bound of Deletion Capacity Sekhari et al. (2021)).
$m_{\epsilon,\delta,\gamma}^{A,M}(d,n)\leq cn$ (18)
A comprehensive proof of Theorem 7 is available in Sekhari et al. (2021). Furthermore, based on Theorem 6, we present the lower bound of deletion capacity below:
Theorem 8 (Deletion Capacity Guarantee).
$m_{\epsilon,\delta,\gamma}^{\Omega,\bar{\Omega}}(d,n)=\mathcal{O}\left(({1}/{\rho})^{n}/{d^{\frac{1}{2}}}\right).$ (19)
Proof of Theorem 8.
Based on Theorem 6, we have
$F(\tilde{\mathbf{w}}_{E-1,B}^{-U})-\mathbb{E}[F\left(\mathbf{w}^{*}\right)]=\mathcal{O}\Big{(}\frac{4L^{2}}{\lambda(n-m)}+\rho^{n{E}/{|\mathcal{B}|}}\frac{2L^{2}}{\lambda}+\rho^{n{E}/{|\mathcal{B}|}}\frac{2L\eta G\sqrt{d}\sqrt{\ln{(1/\delta)}}}{\epsilon}\!\Big{)}.$ (20)
Note that $\rho<1$. The above upper bound on the excess risk thus implies that we can delete at least
$m=\mathcal{O}\Big{(}\frac{\epsilon}{\eta\sqrt{\ln{(1/\delta)}}d^{\frac{1}{2}}\rho^{n}}\Big{)}$ (21)
samples while still ensuring an excess risk guarantee of $\gamma$. Accordingly, it yields an order of $\mathcal{O}\left(({1}/{\rho})^{n}/{d^{\frac{1}{2}}}\right)$. Therefore, we have completed the proof. ∎
B.3 Hessian-Vector Products Technique in Stage II of Algorithm 1
We provide the detailed Hessian-Vector Products (HVP) procedure in Algorithm 2. Notably, when recursively computing the product of the matrix and an arbitrary vector $(\mathbf{I}-{\eta}\mathbf{H})\mathbf{v}$, we can derive it from $\mathbf{v}-{\eta}\mathbf{H}\mathbf{v}$. Here, $\mathbf{H}\mathbf{v}$ involves a HVP through modern automatic differentiation frameworks such as JAX or PyTorch. For more explanation of the HVP, please refer to Dagréou et al. (2024).
1 Stage II: Pre-computing and Pre-storing statistics $\mathcal{T}(\mathcal{S})=\{\mathbf{a}_{E,B}^{-u_{j}}\}_{j=1}^{n}$:
2 initialization $\{\mathbf{a}_{E,B}^{-u_{j}}\}_{j=1}^{n}=0$, $\{\mathbf{v}_{e,b}^{-u_{j}}\}_{j=1}^{n}=0$;
3 for  $e=0,\cdots,E-1,E$ do
4       for  $b=0,\cdots,B-1,B$ do
5             for  $j=1,2\cdots,n$ do
6                   Compute: $\mathbf{a}_{e,b}^{-u_{j}}\leftarrow(\mathbf{I}-{\eta_{e,b}}\mathbf{H}_{e,b})\mathbf{a}_{e,b}^{-u_{j}}+\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}(\mathbf{I}-{\eta_{e,b}}\mathbf{H}_{e,b})\mathbf{v}_{e,b}^{-u_{j}}$
7                   if gradient of sample ${u_{j}}$ computed in this update step then
8                         $\mathbf{v}_{e,b}^{-u_{j}}\leftarrow\nabla\ell(\mathbf{w}_{e,b}^{-u_{j}};z_{i})$
9                  else
10                        $\mathbf{v}_{e,b}^{-u_{j}}\leftarrow 0$
11                   end if
12
13             end for
14
15       end for
16
17 end for
Algorithm 2 Matrix-Vector Products in Stage II of Algorithm 1
B.4 Unlearning-Repairing Strategy in Stage III of Algorithm 1
While our approach does not require accessing any datasets in Stage III, we can mitigate the performance degradation caused by excessive deletions and make our algorithm more robust through a simple fine-tuning strategy if we can still access the remaining dataset in Stage III. We only consider this as a preliminary attempt in this section; other experiments do not include this Finetune strategy. This strategy allows our method to be used as a heuristic algorithm in non-privacy scenarios, without the need for theoretical certified guarantees, such as in bias removal.
Specifically, when the continuous incoming deletion requests to forget data, we can rapidly implement data removal through vector addition. This involves augmenting the current model with approximators, where errors accumulate during this process and thus cause performance degradation. Suppose, upon the arrival of a new round of deletion requests for a particular model and dataset, the model’s accuracy significantly drops due to the removal of a substantial amount of data. As a remedy in Algorithm 3, we can address this performance degradation by fine-tuning the current model for $E_{r}$ epochs on the remaining dataset, referred to as the ‘repairing’ process. Additionally, by computing the approximators for the remaining dataset during the repairing process, and adding these approximators during the repairing process to the pre-stored approximators during the learning process, we obtain the new set of approximators for the remaining data. In Figure 3, we conduct an experiment on ResNet-18 using the CelebA dataset to demonstrate the efficacy of the repairing strategy.
Figure 3: Unlearning-Repairing Strategy. The blue, red, and green lines respectively represent the learning, unlearning, and repairing process. When a significant amount of data is removed, leading to a loss in model performance, performing fine-tuning on the remaining dataset and recording the approximators during this period can help avoid performance degradation.
1 Stage III: Unlearning when user requests to forget the subset $U$=$\{u_{j}\}_{j=1}^{m}$ on model $\mathbf{w}_{E,B}$:
2 Compute: $\bar{\mathbf{w}}_{E,B}^{-U}\!\leftarrow\!\mathbf{w}_{E,B}\!+\!\mathbf{a}_{E,B}^{-U},$ $c\!\leftarrow\!\|\Delta_{E,B}^{-U}\|\frac{\sqrt{2\!\ln\!{(1.25/\delta})}}{\epsilon}\!\!$    Delete statistics: $\{\mathbf{a}_{E,B}^{-u_{j}}\}_{j=1}^{m}$
3 Sample: $\mathbf{n}\sim\mathcal{N}(0,\sigma\mathbf{I})$, Return: $\tilde{\mathbf{w}}_{E,B}^{-U}=\bar{\mathbf{w}}_{E,B}^{-U}+\mathbf{n}$
4Reparing (Fine-tuning) the model $\tilde{\mathbf{w}}^{-U}_{E,B}$ on the remaining dataset $S\backslash U=\{z_{r}\}_{r=1}^{n-m}$:
5 for $e=0,...,E_{r}$ do
@@ B.4 Unlearning-Repairing Strategy in Stage III of Algorithm 1
7             Compute gradient: $\mathbf{g}\!\leftarrow\!\frac{1}{|\mathcal{B}_{e,b}|}\sum_{r\in\mathcal{B}_{e,b}}\!\!\!\nabla\ell\left(\mathbf{w}_{e,b};z_{r}\right)$,
@@ 8             Gradient descent: $\mathbf{w}_{e,b+1}\leftarrow\mathbf{w}_{e,b}-{\eta_{e,b}}{\mathbf{g}}$
12
13Pre-computing statistics $\mathcal{T}(\mathcal{S})=\{\mathbf{a}_{E,B}^{-u_{r}}\}_{r=1}^{n-m}$ of the remaining dataset $S\backslash U=\{z_{r}\}_{r=1}^{n-m}$:
14 for  $r=1,2\cdots,{n-m}$ do
15       Recursive computation by using HVP based on Algorithm 2:
16       $\mathbf{a}_{E+E_{r},B}^{-u_{r}}\leftarrow\sum_{e=0}^{E_{r}}\mathbf{M}_{e,b(u_{r})}\cdot\nabla\ell(\mathbf{w}_{e,b(u_{r}))};u)+\hat{\mathbf{H}}_{E_{r},B-1\rightarrow 0,b(u_{r})+1}\cdot\mathbf{a}_{E,B}^{-u_{r}}$
17       where: $\mathbf{M}_{e,b(u_{r})}=\frac{\eta_{e,b(u_{r})}}{|\mathcal{B}_{e,b(u_{r})}|}\hat{\mathbf{H}}_{E,B-1\rightarrow e,b(u_{r})+1}$
@@ 15       Recursive computation by using HVP based on Algorithm 2:
19
Algorithm 3 Hessian-Free Unlearning-Repairing Strategy in Stage III of Algorithm 1
Appendix C Proofs
C.1 Detailed Analysis of Eq. 11
Before proceeding with the analysis of our method, we will first provide the analysis of previous works (Suriyakumar et al. (2022); Liu et al. (2023); Mahadevan & Mathioudakis (2021); Chien et al. (2022); Liu et al. (2023)) and discuss the circumstances that lead to its failure.
We first assume that the empirical risk in Eq. 2 is twice-differentiable and strictly convex in the parameter space $\mathcal{W}$. The original empirical risk minimizer can be expressed as
$\hat{\mathbf{w}}:=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{W}}\quad F_{\mathcal{S}}=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{W}}\quad\frac{1}{n}\sum_{i=1}^{n}\ell(\mathbf{w};z_{i}).$ (22)
We reweight the sample $u$ in the original empirical risk with a weighting factor $\omega$,
$\hat{\mathbf{w}}^{-u}:=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{W}}$ $\quad\frac{1}{n}\sum_{i=1}^{n}\ell(\mathbf{w};z_{i})+\omega\ell(\mathbf{w};u),$ (23)
$\text{where}\quad\omega=-\frac{1}{n}.$
When weighting factor $\omega=-\frac{1}{n}$, Eq. 23 is equivalent to the empirical risk minimizer without samle $u$, i.e. removing a training point $u$ is similar to up-weighting its corresponding weight by $\omega=-\frac{1}{n}$. Following the classical result of Weisberg & Cook (1982), Koh & Liang (2017) aimed to approximate $\hat{\mathbf{w}}^{-u}$ by the first-order Taylor expansion around the optimal learned model $\hat{\mathbf{w}}$, given by,
$\hat{\mathbf{w}}^{-u}-\hat{\mathbf{w}}\approx-\frac{1}{n}\left.\frac{d\hat{\mathbf{w}}^{-u}}{d\omega}\right|_{\omega=0}=\left(\frac{1}{n}\sum_{i=1}^{n}\nabla^{2}\ell(\hat{\mathbf{w}};z_{i})\right)^{-1}\nabla\ell(\hat{\mathbf{w}};u).$ (24)
However, we note that these works rely on several key assumptions: the convexity of the objective function to ensure the invertibility of the Hessian matrix, and the optimality of the learned model. These assumptions are often stringent and may not be satisfied in practice, particularly in modern machine learning scenarios where the objective function may be non-convex and the empirical risk minimizer may not be attainable. In addition, in over-parameterized models, the computation of the Hessian matrix also requires a significant memory overhead.
In fact, aside from the optimal model, the training process contains more information, and our method aims to utilize the information from the training process to address both issues simultaneously. Here, we provide a more detailed explanation for Eq. 11. Specifically, we first consider the impact of sample $u$ during the $e$-th epoch, and then extend the analysis to all epochs. It should be noted that a necessary setting for our method is that the retraining and learning process have the same initialization, model architecture, and training process.
Analysis of Eq. 11.
For ease of understanding, each sample appears only in one (mini-)batch per epoch. The update rule for the model parameters is typically given by,
$\mathbf{w}_{E,B+1}\leftarrow\mathbf{w}_{E,B}-\frac{\eta_{E,B}}{|\mathcal{B}_{E,B}|}\sum_{i\in\mathcal{B}_{E,B}}\nabla\ell\left(\mathbf{w}_{E,B};z_{i}\right).$ (25)
Without loss of generality, for the $e$-th epoch, we define the sample $u$ to be sampled in the $e$-th epoch’s $b(u)$-th update. Reweighting the sample $u$ in the each update process (rather than the empirical risk) with a weighting factor $\omega$, Eq. 3 and Eq. 4 can be represented by